# Effective-medium approach to the resonance distribution of wave scattering in a random point field

### v2

@@ IV.2 Semiclassical transport equation
Therefore, we find reasonable to neglect these effects in this section.
Since only the density appears in the right-hand side of Eq. (68), this equation is closed for the diagonal part of the density matrix [28], which is nothing but the probability density $\langle\rho(\boldsymbol{\mathrm{r}},\boldsymbol{\mathrm{r}})\rangle=\langle\rho(\boldsymbol{\mathrm{r}})\rangle=\langle\left|\psi(\boldsymbol{\mathrm{r}})\right|^{2}\rangle$. The fact that the equation is closed for $\langle\rho(\boldsymbol{\mathrm{r}})\rangle$ results from the assumption that the scatterers are point-like in the Foldy-Lax model. As a consequence, the scattering amplitude $F(k)$ is isotropic, and represented by a scalar number, instead of a momentum-dependent operator. Thus, on each collision, the particle loses the memory of its initial direction, and can be represented on each point by an isotropic spherical wave. Therefore, the only relevant distribution to describe the particle state is the density $\langle\rho(\boldsymbol{\mathrm{r}})\rangle$. However, in the more general case of directional scattering, the equation for the density matrix $\langle\rho(\boldsymbol{\mathrm{r}},\boldsymbol{\mathrm{r}}^{\prime})\rangle$ could not be simplified to an equation for $\langle\rho(\boldsymbol{\mathrm{r}})\rangle$.
In the isotropic scattering case, it is convenient to define a new integral kernel with the square modulus of the effective Green function
$K(\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})\xrightarrow{r\rightarrow\infty}\frac{\mathop{}\!\mathrm{e}^{-(\gamma+n\sigma)\left\|\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}^{\prime}\right\|}}{S_{d}\left\|\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}^{\prime}\right\|^{d-1}}\>,$ (71)
where $\gamma=\tfrac{s}{v}=2\operatorname{Im}k$ is the dual variable of time under Laplace transform.
Using the kernel (70) and the point cross section formula (14), the diagonal part of the Bethe-Salpeter equation (68) reads
$\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle=K(\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}_{0})+\int_{\mathbb{R}^{d}}K(\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})n(\boldsymbol{\mathrm{r}}^{\prime})\sigma(k)\left\langle\rho(\boldsymbol{\mathrm{r}}^{\prime})\right\rangle\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{r}}^{\prime}\>.$ (72)
Note that the small imaginary part of $k=k_{0}+\mathrm{i}{k}_{\rm i}$ can be neglected in the cross section because the equation only considers time scales much longer than the oscillation period of the wave ($s\ll\omega_{0}$ and ${k}_{\rm i}\ll k_{0}$). It is thus safe to write $\sigma(k)\simeq\sigma(k_{0})$. Note also that, from Eq. (68) to Eq. (72), we have changed the normalization of the density because the source term $\left|\langle Q(k,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}_{0})\rangle\right|^{2}$ became $K(\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}_{0})$ defined in Eq. (70). However, this has no consequence on $\langle\rho(\boldsymbol{\mathrm{r}})\rangle$ except for an irrelevant global factor.
The formal solution of the transport equation (72) reads
@@
In Sec. IV.5, we will study the spectrum of the solution (73) in the variable $\gamma$ introduced in Eq. (45), so as to make the connection with the band of complex resonances at $\left|k\right|\varsigma\gtrsim j_{\frac{d-2}{2}}$. Indeed, according to the relationship (64) between the density matrix and the resolvent, we expect that this solution predicts the depth of the resonance band in the complex plane of $k$.
Furthermore, the transport equation (72) is equivalent to the linear Boltzmann equation for the Wigner function of the quantum particle [25, 33, 34], as shown in Appendix G. This reveals an interesting connection with decoherence theory of a quantum particle in a gas of moving atoms [74, 75, 76, 77, 78]. In the Foldy-Lax model, where the scatterers are fixed and the disorder is quenched, the randomness of the scatterer positions leads to decoherence, which is characterized by a finite coherence length [79, 80, 77, 78]
$\Lambda(\boldsymbol{\mathrm{r}}_{0})^{2}=\frac{\int_{\mathbb{R}^{d}}\left\|\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}_{0}\right\|^{2}\left|\rho(\boldsymbol{\mathrm{r}},\boldsymbol{\mathrm{r}}_{0})\right|^{2}\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{r}}}{\int_{\mathbb{R}^{d}}\left|\rho(\boldsymbol{\mathrm{r}},\boldsymbol{\mathrm{r}}_{0})\right|^{2}\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{r}}}\>.$ (74)
The length (74) is nothing but the standard deviation of the distribution $\left|\rho(\boldsymbol{\mathrm{r}},\boldsymbol{\mathrm{r}}^{\prime})\right|^{2}$ around a reference point, say $\boldsymbol{\mathrm{r}}^{\prime}$. This is a measure of the characteristic length beyond which the statistical correlations are lost. Indeed, the average density matrix is also the autocorrelation function of the wavefunction, i.e., $\langle\rho(\boldsymbol{\mathrm{r}},\boldsymbol{\mathrm{r}}^{\prime})\rangle=\langle\psi(\boldsymbol{\mathrm{r}}){\psi}^{*}(\boldsymbol{\mathrm{r}}^{\prime})\rangle$. Beyond the coherence length, the wave formalism becomes less meaningful in the sense that two wave packets located far enough from each other are statistically independent and experience different collision histories without possible interference between them.
Let us determine the coherence length (74) in the special case of an infinite disordered medium and at equilibrium, that is after a time much longer than the mean intercollisional time. It is shown in Appendix H that, in this regime, the density matrix obeying Eq. (68) asymptotically behaves as
$\begin{split}\rho_{\infty}(\boldsymbol{\mathrm{r}},\boldsymbol{\mathrm{r}}^{\prime})\xrightarrow{r\rightarrow\infty}&\frac{1}{2k_{0}}\left(\frac{k_{0}}{2\pi r}\right)^{\frac{d-1}{2}}\\ &\times\sin\left(k_{0}r-\tfrac{\pi(d-3)}{4}\right)\mathop{}\!\mathrm{e}^{-\frac{n\sigma}{2}r}\>,\end{split}$ (75)
where $r=\left\|\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}^{\prime}\right\|$. Substituting Eq. (75) into Eq. (74) provides the sought coherence length
$\Lambda(\boldsymbol{\mathrm{r}}_{0})=\sqrt{2}\,\ell_{\rm s}\>.$ (76)
This result shows that the coherence length is of the order of the mean free path. In this model, the mean free path is thus not only the mean distance between two successive collisions, but also the coherence length. Due to the analogy with decoherence theory [74, 75, 76, 81, 82, 77, 78] mentioned before, the result (76) confirms the interpretation of the exponential evanescence of the average Green function at the scale of the mean free path highlighted in Eq. (46) as being related to decoherence.
IV.3 Diffusion approximation
The issue with the transport equation (72) is that solving it analytically is rarely feasible, even in simple geometries such as the spherical geometry considered in this paper. In addition, this equation is quite difficult to solve numerically because it generally requires too large discretization lattices to be practical. However, it is possible to considerably simplify the problem in the diffusion regime, that is when the mean free path $\ell_{\rm s}$ is much smaller than the radius $R$ of the spherical medium.
In this section, we compare the average density $\langle\rho(\boldsymbol{\mathrm{r}})\rangle$ predicted in the diffusion approximation to the actual density of the wavefunction for both a single realization of the disorder and on average. In addition, we show that the diffusion approximation leads to an accurate estimate of the depth of the resonance band in Figs. 3(a) and 3(b) with respect to the real $k$ axis for $d\geq 2$.
As shown in Appendix I, Eq. (72) reduces in the diffusion regime and far away from the boundaries to
$\gamma\left\langle\rho(\boldsymbol{\mathrm{r}},\gamma)\right\rangle=\delta(\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}_{0})+\frac{1}{dn\sigma}\nabla^{2}_{\boldsymbol{\mathrm{r}}}\left\langle\rho(\boldsymbol{\mathrm{r}},\gamma)\right\rangle\>.$ (78)
To let appear the diffusion equation out of Eq. (78), we have to perform the inverse Laplace transform in time
$\left\langle\rho(\boldsymbol{\mathrm{r}},t)\right\rangle=\frac{1}{2\pi\mathrm{i}}\int_{-\mathrm{i}\infty}^{\mathrm{i}\infty}\mathop{}\!\mathrm{e}^{\gamma vt}\left\langle\rho(\boldsymbol{\mathrm{r}},\gamma)\right\rangle\mathop{}\!\mathrm{d}(\gamma v)\>,$ (79)
recalling that $s=\gamma v$. This is why, in Eq. (78), we have stressed the dependence of $\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle$ on $\gamma$. We ultimately obtain the diffusion equation
$\frac{\partial}{\partial t}\left\langle\rho(\boldsymbol{\mathrm{r}},t)\right\rangle-D\nabla^{2}_{\boldsymbol{\mathrm{r}}}\left\langle\rho(\boldsymbol{\mathrm{r}},t)\right\rangle=\delta(\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}_{0})\delta(t)\>,$ (80)
where the parameter $D$ is the diffusivity, or diffusion constant, characterizing the diffusive transport of the particle in the medium. It reads
$D=\frac{v}{dn\sigma}=\frac{v\ell_{\rm s}}{d}\>.$ (81)
Note that the right-hand side of Eq. (80) contains a source term at $\boldsymbol{\mathrm{r}}=\boldsymbol{\mathrm{r}}_{0}$ at time $t=0$. This is consistent with the assumptions made in Sec. IV.1 about the initial wave packet $\hat{\rho}_{0}=\left|\boldsymbol{\mathrm{r}}_{0}\right\rangle\left\langle\boldsymbol{\mathrm{r}}_{0}\right|$. The diffusion equation (80) can be solved in an infinite medium. Setting $\boldsymbol{\mathrm{r}}_{0}=\boldsymbol{\mathrm{0}}$, the result is [25]
$\left\langle\rho(\boldsymbol{\mathrm{r}},t)\right\rangle=\frac{1}{(4\pi Dt)^{\frac{d}{2}}}\exp\left(-\frac{r^{2}}{4Dt}\right)\Theta(t)\>,$ (82)
where $\Theta(t)$ denotes the Heaviside step function. Expression (82) is a normalized Gaussian distribution. The variance of this distribution is given by
$\left\langle\boldsymbol{\mathrm{r}}(t)^{2}\right\rangle=\int_{\mathbb{R}^{d}}\left\|\boldsymbol{\mathrm{r}}\right\|^{2}\left\langle\rho(\boldsymbol{\mathrm{r}},t)\right\rangle\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{r}}=2dDt=2\ell_{\rm s}vt\>,$ (83)
and thus linearly increases in time.
The diffusion equation must be supplemented by the condition on the spherical boundary ($\forall\boldsymbol{\mathrm{r}}\in\mathcal{S}_{d}(R)$),
$\left(\gamma+\frac{n\sigma}{2}\right)\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle+\frac{V_{d-1}}{S_{d}}\frac{n\sigma}{\gamma+n\sigma}\boldsymbol{\mathrm{n}}\cdot\boldsymbol{\mathrm{\nabla}}_{\boldsymbol{\mathrm{r}}}\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle=0\>,$ (84)
which is derived at the end of Appendix I. When considering times much longer than the intercollisional time, the variable $\gamma$ in Eq. (84) can be neglected in front of $n\sigma$. In this way, we obtain a time-independent version of Eq. (84), or rather a $\gamma$-independent version:
$\frac{n\sigma}{2}\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle+\frac{V_{d-1}}{S_{d}}\boldsymbol{\mathrm{n}}\cdot\boldsymbol{\mathrm{\nabla}}_{\boldsymbol{\mathrm{r}}}\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle=0\>.$ (85)
This is a Robin boundary condition on the density $\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle$ for all $\boldsymbol{\mathrm{r}}\in\mathcal{S}_{d}(R)$. Since the coefficients of the two terms in Eq. (85) are positive, this condition expresses the fact that the density has an inward-facing gradient. In the far diffusive regime ($n\sigma R\rightarrow\infty$), this condition reduces to a simple Dirichlet boundary condition: $\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle=0$. Although this regime is valid for a very large medium, we do not consider this extreme case in this paper.
IV.4 Spatial profile of the wavefunction
In this section, we compare the predictions of the diffusion equation regarding the average density $\langle\rho(\boldsymbol{\mathrm{r}})\rangle$ to the numerical results obtained in the Foldy-Lax model with the program MSModel [56]. In particular, we want to know if the disordered-average density, $\langle\left|\psi(\boldsymbol{\mathrm{r}})\right|^{2}\rangle$, is representative of the density in a single realization of the disorder. In other words, we wonder if the system has some self-averaging property.
Furthermore, we want to compare the prediction of the diffusion equation (78) to the numerical value of $\langle\left|\psi(\boldsymbol{\mathrm{r}})\right|^{2}\rangle$ in the spherical-shaped point field. For this purpose, we assume that the wave packet is created at the center of the spherical medium ($\boldsymbol{\mathrm{r}}_{0}=\boldsymbol{\mathrm{0}}$). In the stationary regime ($\gamma\rightarrow 0$), the Laplace-transformed diffusion equation (78) reads
$\nabla^{2}_{\boldsymbol{\mathrm{r}}}\rho_{\rm da}(\boldsymbol{\mathrm{r}})=-A\,\delta(\boldsymbol{\mathrm{r}})\>,$ (86)
where the subscript stands for diffusion approximation, and the amplitude factor is $A=dn\sigma$. However, the normalization convention in our numerical simulations is slightly different from what Eq. (78) prescribes. In these simulations, we assume that the source term in Eq. (72) is $\left|\langle Q(k,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}_{0})\rangle\right|^{2}$ and not $K(\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}_{0})$. The correction is simply given by the prefactor in Eq. (70). Thus, we should have
$A=dn\sigma\frac{\pi\mu(k_{0})}{k_{0}}\>.$ (87)
The general solution of the Poisson equation (86) reads
$\rho_{\rm da}(r)=\begin{cases}-\frac{A}{S_{d}}\frac{r^{2-d}-R_{\rm eff}^{2-d}}{2-d}&\text{for}~{}d\neq 2\>,\\[10.0pt] -\frac{A}{2\pi}\ln\left(\frac{r}{R_{\rm eff}}\right)&\text{for}~{}d=2\>,\end{cases}$ (88)
where the arbitrary parameter $R_{\rm eff}$ denotes the effective radius at which the density is supposed to vanish as if the medium was extended beyond $R$. Since the density must be positive everywhere, the effective radius is necessarily larger or equal to the actual radius $R$ of the medium ($R_{\rm eff}\geq R$). The effective radius has to be adjusted to satisfy the boundary condition (85), which can also be rewritten as
$\frac{\rho_{\rm da}(R)}{\partial_{r}\rho_{\rm da}(R)}=-\frac{2V_{d-1}}{S_{d}}\ell_{\rm s}\>.$ (89)
In order to satisfy Eq. (89), the effective radius $R_{\rm eff}$ in Eq. (88) must read
$R_{\rm eff}=\begin{cases}R\left(1+(2-d)\frac{2V_{d-1}}{S_{d}}\frac{\ell_{\rm s}}{R}\right)^{\frac{1}{2-d}}&\text{for}~{}d\neq 2\>,\\[10.0pt] R\exp\left(\frac{2V_{d-1}}{S_{d}}\frac{\ell_{\rm s}}{R}\right)&\text{for}~{}d=2\>,\end{cases}$ (90)
Note that, if the mean free path is small enough ($\ell_{\rm s}\ll R$), the expressions (90) can be effectively expanded at first order. We find
$R_{\rm eff}\simeq R+\frac{2V_{d-1}}{S_{d}}\ell_{\rm s}\>.$ (91)
Since the dimensionless factor $2V_{d-1}/S_{d}$ is of the order of one, the result (91) means that the effective radius is about the actual radius plus the mean free path. This is a known result of transport theory, which is, for instance, encountered in the neutronics of nuclear reactors [34].
The square modulus of the wavefunction numerically computed with the program MSModel [56] in a spherical-shaped random point field for $N=10^{3}$ is shown in Fig. 4 and the cross-sectional view long the $x$ axis in Fig. 5. The similar cross-sectional view of Fig. 1(a) is shown in Fig. 6.
Figure 4: (a) Square modulus of the wavefunction computed with the program MSModel [56] in a 2D disk-shaped random point field of $N=10^{3}$ scatterers with $k=10\,\varsigma^{-1}$. The point scattering model is Eq. (15) and the mean free path is $\ell_{\rm s}=2.5\,\varsigma$. The point source is located at the center ($\boldsymbol{\mathrm{r}}_{0}=\boldsymbol{\mathrm{0}}$). The circle depicts the medium boundary. (b) Same as panel (a) but averaged over $2^{6}$ random configurations.
Figure 5: (a) Cross-sectional view of Fig. 4(a) along the positive $x$ semiaxis. The vertical gray line indicates the radius of the medium. The dashed curve is the diffusion approximation (88) using Eqs. (87) and (90). (b) Cross-sectional view of Fig. 4(b) averaged over $2^{14}$ configurations. The filled region is the numerically computed interquartile range of $\left|\psi(\boldsymbol{\mathrm{r}})\right|^{2}$. The dotted curve is the square modulus of the effective Green function (40).
Figure 6: Cross-sectional view of Fig. 1(a) along the positive $x$ semiaxis. The vertical gray line indicates the radius of the medium. The dashed curve is the diffusion approximation (88) using Eqs. (87) and (90).
We note that the filament-like structures observed in Fig. 1(a) and Fig. 4(a) have a scale of the order of the mean free path $\ell_{\rm s}$ introduced by Eq. (44) in relation to the free flights of the particle between its collisions on the scatterers.
Figure 4(b) shows the wavefunction density averaged over several random configurations of the scatterers. We observe that the detailed structures of the wavefunction seen in Fig. 4(a) and, in particular, the random meandering nodal lines have disappeared after averaging.
In both Figs. 5 and 6, the dashed curve is the theoretical average density (88) predicted by the diffusion approximation. The matching with the numerical average for $r\leq R$ in Fig. 5(b) is remarkable, despite the apparent roughness of the diffusion approximation. Indeed, the scale parameter $R/\ell_{\rm s}\simeq 7$ is not so large compared to one. Figures 5(a) and 6 seem to confirm that the wavefunction in a single configuration globally follows the theoretical decrease of $\rho_{\rm da}(r)$. This suggests that the aspect of the wave density in an infinite disordered medium will be closer to Fig. 4(b) than to Fig. 4(a), because the medium will ultimately look uniform at such large scales. This would support the validity of the disorder averaging for a sufficiently large medium ($R\gg\ell_{\rm s}$).
In the neighborhood of the source point at $\boldsymbol{\mathrm{r}}_{0}=\boldsymbol{\mathrm{0}}$, the numerical average does not match the diffusion approximation in Fig. 5(b). The reason is that the diffusion approximation is not supposed to be accurate on scales shorter than the mean free path. Indeed, the amplitude of the singularity predicted by the diffusion approximation is not consistent with that of the Bethe-Salpeter equation (72). According to Eq. (68), the intensity near the source point $\boldsymbol{\mathrm{r}}_{0}$ is approximately given by the effective Green function $\left|\langle Q(k,\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}_{0})\rangle\right|^{2}$. This is confirmed by the matching between the solid and the dotted curves near $r=0$ in Fig. 5(b).
Besides that, one notices a mismatch at the edge of the medium for $r\geq R$. This is due to the fact that the density (88) is only valid within the point field. In the outer region ($r\geq R$), the density behaves as the power law
$\rho(r)\propto\frac{1}{r^{d-1}}\>.$ (92)
This is given by the far-field behavior of Eq. (72) for $r\gg R$.
IV.5 Depth of the resonance band
The integral equation (72) also provides a prediction of the position of the band of complex resonances seen in Figs. 3(a) and 3(b) but also in our previous paper [47]. The resonant states can be obtained as the eigenvectors of the Schrödinger equation without any source term, just like eigenstates but at complex energies. In the same way, in order to study the resonance band from Eq. (72), we need to set the source term $K(\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}_{0})$ to zero and interpret the resulting equation
@@ IV.5 Depth of the resonance band
$\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle=\int_{\mathbb{R}^{d}}\frac{\mathop{}\!\mathrm{e}^{-(\gamma+n\sigma)\left\|\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}^{\prime}\right\|}}{S_{d}\left\|\boldsymbol{\mathrm{r}}-\boldsymbol{\mathrm{r}}^{\prime}\right\|^{d-1}}n(\boldsymbol{\mathrm{r}}^{\prime})\sigma(k)\left\langle\rho(\boldsymbol{\mathrm{r}}^{\prime})\right\rangle\mathop{}\!\mathrm{d}\boldsymbol{\mathrm{r}}^{\prime}\>,$ (93)
as an eigenvalue problem, where the free variable $\gamma$ plays the role of the sought eigenvalue and the density $\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle$ is the associated eigenfunction. The eigensystem (93) is more easily solved in the diffusion approximation. Using Eqs. (78) and (85), the integral equation (93) can be cast into

### v3

@@ II.4 Semiclassical transport
Therefore, we find reasonable to neglect these effects in this paper.
The formal solution of the transport equation (39) is given by
@@
As we will see in Sec. III, the integral equation (39) provides a prediction of the position of the peaks in the resonance density observed in our previous paper [47]. According to Eqs. (38) and (43), we expect that most of the complex resonances occur in the vicinity of the singularities of $\langle\rho(\boldsymbol{\mathrm{r}},\gamma)\rangle$ in the variable $\gamma$. These singularities are given by the eigenvalue problem
@@
$\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle=\int_{\mathbb{R}^{d}}\mathop{}\!\mathrm{d}{\boldsymbol{\mathrm{r}}^{\prime}}K(\boldsymbol{\mathrm{r}}\mid\boldsymbol{\mathrm{r}}^{\prime})n(\boldsymbol{\mathrm{r}}^{\prime})\sigma(k)\left\langle\rho(\boldsymbol{\mathrm{r}}^{\prime})\right\rangle\>,$ (44)
where the free variable $\gamma$ (defined in Eq. (36) and used in Eq. (41)) plays the role of the sought eigenvalue and the density $\left\langle\rho(\boldsymbol{\mathrm{r}})\right\rangle$ is the associated eigenfunction. The singular values $\gamma$ given by Eq. (44) are expected to correspond to the main structures observed in the resonance density $\varrho^{(2)}({k}_{\rm r},{k}_{\rm i})$ in the high-energy regime where the coherent (phase-dependent) effects are negligible.
III Numerical results
In this section, the predictions of the effective-medium equations of Secs. II.3 and II.4 are compared to the exact density of complex resonances obtained numerically using the resonance potential method developed in our previous paper [47]. In that paper, two distinctive structures are observed in the complex resonance spectrum:
  1. 1.
    Peaks at low energy, that is in the region
    $\left|k\right|\varsigma\lesssim j_{\frac{d-2}{2}}\>,$ (45)
    where $\varsigma$ is the mean interscatterer distance (7) and $j_{\frac{d-2}{2}}$ is the first zero of the Bessel function $J_{\frac{d-2}{2}}(z)$. These peaks are explained in Sec. III.1 by the Dyson-type effective-medium equation (30).
  2. 2.
    A nearly horizontal resonance band at larger energies which is characterized in Sec. III.2 using the Bethe-Salpeter equation of the form (44).
III.1 Low-energy peaks in the resonance density
We first consider the low-energy peaks in the resonance density shown in Fig. 2.
Figure 2: (a) Resonance density (28) computed with the program MSModel [56] in a 2D disk-shaped random point field for $N=50$ and using the point scattering model (15) with $\alpha=0.1\,\varsigma$. The parameters are thus $R=3.99\,\varsigma$ and $\ell_{\rm s}\simeq 1\,\varsigma$ (multiple-scattering regime). The density is averaged over $2^{8}$ disorder configurations. The resonances of the effective medium, obtained by numerically solving Eq. (46), are depicted by small circles labelled with the corresponding angular momentum $\ell$. The dashed line demarcates the low-energy region (45). (b) Same as panel (a) but in a 3D spherical random point field for $N=100$. The parameters are thus $R=2.88\,\varsigma$ and $\ell_{\rm s}\simeq 10\,\varsigma$ (quasiballistic regime).
We solve the effective-medium equation (30) for the resonance poles in a spherical disordered medium of radius $R$. The details of this calculation are given in Appendix A using the partial wave method [57, 58, 59, 60], and the result is the equation
for a complex resonance at $k\in\mathbb{C}$. In Eq. (46), $\kappa$ is the effective wavenumber (31) with $n(\boldsymbol{\mathrm{r}})=n$, $J_{\nu}(z)$ and $H^{+}_{\nu}(z)$ are the Bessel and Hankel functions, respectively, and the order $\nu$ is given by
$\nu=\ell+\frac{d-2}{2}\>,$ (47)
where $\ell$ is the orbital quantum number ($\ell\in\{0,1,2,\ldots\}$).
The comparison between the numerical solutions of Eq. (46) for $k\in\mathbb{C}$ and the actual peaks of the resonance density obtained is shown in Fig. 2(a) for a 2D point field and in Fig. 2(b) for a 3D point field. In these two panels, instead of Eq. (16), we use the point scattering model (15) with the scattering length $\alpha=0.1\,\varsigma$ to avoid the divergent behavior of the maximum cross section in the small-$k$ regime.
The adequacy between the roots of Eq. (46) and the peaks in Figs. 2(a) and 2(b) is remarkable and supports the validity of the effective-medium equation (30) to describe the resonance density at large wavelength ($k\varsigma\ll 1$). It shows that, in this low-energy regime, the incident wave mainly perceives the random point field as a semitransparent sphere without the details due to disorder.
Moreover, it turns out that the resonance structure shown in Fig. 2 is little affected by the geometric shape of the disordered medium or the scattering parameters such as the cross section of the point scatterers. Indeed, the same kind of structure can be observed for a cubic medium, instead of a spherical medium. The robustness of this resonance structure allows us to make a quite radical simplification of the resonance equation (46). Indeed, given the weak influence of the cross section of the point scatterers, it would be possible to let the parameter $\kappa$ tend to infinity, leading to
$H^{+}_{\nu}(kR)=0\>.$ (48)
This is the equation of resonances corresponding to the scattering by a hard sphere of radius $R$. The roots of Eq. (48) are shown in Fig. 3 and indeed form structures similar to those in Fig. 2.
Figure 3: First complex zeros of the Hankel function $H^{+}_{\nu}(z)$ for $\arg z\in[-\pi,0]$ at integer and half-integer orders (even and odd dimensions respectively). The dotted curves represent the trajectory of the zeros when $\nu$ is continuously increased. Zeros at the same order are connected by a solid line. The other zeros around $\arg z=-\pi$ for $\nu\in\mathbb{Z}$ are not shown.
This observation confirms the robustness of this low-energy resonance structure against possibly significant changes of the scattering parameters. Nevertheless, the hard-sphere resonance equation (48) is not as accurate as Eq. (46) because it completely neglects the scattering within the medium. This is why the roots of Eq. (48) are not shown in Fig. 2. In contrast, the prediction of the resonance equation (46) of the effective-medium theory is more reliable because it includes the effects of finite values of $\kappa$.
In the 2D case [Fig. 2(a)], a slight shift between simulation and prediction can be noticed. In addition, the predictions in the upper right side of Fig. 2(a) do not match the resonance band close to the real $k$ axis. These disagreements are likely due to multiple-scattering effects because the regime here is diffusive ($R/\ell_{\rm s}\simeq 4$). Indeed, this would explain why the agreement with Eq. (46) is better in the 3D case [Fig. 2(b)] where the scattering regime is quasiballistic ($R/\ell_{\rm s}\simeq 0.3$). In order to take into account multiple-scattering effects, we need to go beyond the approach used in this section, which is based solely on the effective-medium equation (30). One possible strategy would be to solve a coupled system formed by Eqs. (30), (40), and (44) for singular values of $k\in\mathbb{C}$.
III.2 Depth of the resonance band
Now, we consider the band of resonance poles almost parallel to the real $k$ axis and visible in the upper right corner of Figs. 2(a) and 2(b). In this high-energy region where $\left|k\right|\varsigma\gtrsim j_{\frac{d-2}{2}}$, we assume that coherent effects are negligible and that the transport is semiclassical so that we may use the Bethe-Salpeter equation of the form (44). In order to solve this eigenproblem, we use the diffusion approximation which holds when the mean free path is much smaller than the characteristic size of the disordered medium ($\ell_{\rm s}\ll R$). As shown in Appendix B, the integral equation (44) can be cast in this regime into the following diffusion equation and Robin boundary condition: